OpenStudy (deercult):

What is the area of a sector with a central angle of 160 degrees and a diameter of 5.8 m? Round the answer to the nearest tenth.

OpenStudy (freckles):

\[\frac{\text{ Area of sector }}{\text{ Area of circle }}=\frac{\text{central angle }}{ 360^\circ }\]

OpenStudy (freckles):

solve for Area of sector remember area of a circle (you know the full circle ) is pi*r^2

OpenStudy (freckles):

\[\frac{\text{ Area of sector }}{\pi r^2}=\frac{\text{central angle }}{ 360^\circ }\]

OpenStudy (freckles):

multiply both sides by pi*r^2 to get the formula for area of sector

OpenStudy (freckles):

\[\text{ Area of sector }= \pi r^2 \cdot \frac{\text{ central angle}}{360^\circ}\]

OpenStudy (freckles):

this formula is when the central angle is in degrees

OpenStudy (freckles):

which it is here

OpenStudy (freckles):

pi*r^2 should be 8.41*pi not just 8.41

OpenStudy (deercult):

\[360x = 1345.6\pi\] ?

OpenStudy (freckles):

which did you multiply both sides by 360

OpenStudy (freckles):

you are trying to solve for x not 360x :p

OpenStudy (freckles):

\[x=\frac{1345.6 \pi}{360}\]

OpenStudy (freckles):

why did you multiply both sides by 360 *

OpenStudy (deercult):

I was writing down, like, the v first step of the equation. I tend to take things in baby steps, hah

OpenStudy (deercult):

Hm... I think I got it wrong somewhere cause 3.7 isn't exactly a viable answer

OpenStudy (deercult):

Hang on

OpenStudy (freckles):

yea 3.7 isn't correct

OpenStudy (freckles):

looks like you keep ignoring the pi in your calculations

OpenStudy (deercult):

Whoops

OpenStudy (freckles):

did you get your answer yet?

OpenStudy (freckles):

use 3.14 for pi you have 3.73 * 3.14 is the last thing you should have now

OpenStudy (freckles):

i'm sorry i must go

OpenStudy (deercult):

Oh, alright! I'll try and work it out from here.

OpenStudy (deercult):

Got it! 11.7 m^2